Q1: Use Laplace transforms to solve the initial value problem x – 2x + 5x = 1, 2(0) = 0, x(0) = 0. 111. 7177J71111
Answer #1

20 - 2 x! +5% = 1 L&x - 2x + 5x 4 = Lşi 1824 - 22{x4 + 5L{x} = at sxs) - SXD) - 2010) – 2sX 5) + XD) +5XL) = 1 lis25 +5 )

Note: to find the solution of given initial value problem we use lapalce transform method.

Details explained in the image.